Circular No. 6889

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bo IN ESO 185-G31 J. Maza, Department of Astronomy, University of Chile, reports the

discovery by Marina Wischnjewsky, on a T-Max 400 film taken by L. E. Gonzalez with the Maksutov telescope at Cerro El Roble on Apr. 22.290 UT, of an apparent supernova (B ∼ 18) located at α = 19^{h}57^{m}22^{s}.55, δ = −55^{o}08′18′′.4 (equinox 2000.0), which is 3′′.0 east and 5′′.1 south of the host galaxy’s center. SN 1998bo was confirmed on direct CCD images obtained by R. Covarrubias on Apr. 28.338 using the 0.91-m telescope at Cerro Tololo. The object is not present on a film obtained on Mar. 31.4 (B > 20).

F. Patat, European Southern Observatory (ESO); and M. Maia, Observatorio Nacional, Rio de Janeiro, report observations of SN 1998bo with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph; 20-min exposure, range 360–750 nm, resolution 0.5 nm) on Apr. 29.32 UT at La Silla: “A full reduction of the CCD spectrogram shows that this object is indeed a supernova. The spectral features are consistent with a type-Ic object, 2–3 weeks after maximum light, and are quite similar to those shown by the well-studied SN 1987M (Filippenko et al. 1990, A.J. 100, 1575). The most prominent features arise from Ca II, Fe II, Na I, and Si II. The expansion velocities deduced from the minima of Ca II H and K, Fe II 501.8-nm, Fe II 516.9-nm, and Na I D are 14 900, 9400, 8900, and 10000 km/s, respectively. The recession velocity, deduced from the unresolved Hα line arising in the parent galaxy, is 4830 km/s, and this is consistent with the value of 4793 km/s reported by Mathewson (1996, Ap.J. Suppl. 107, 97).”

SUPERNOVA 1998bn IN NGC 4462 Visual magnitude estimate by J. Carvajal, Madrid, Spain: Apr. 28.97

UT, 13.9.

NOVA SAGITTARII 1998 Visual magnitude estimates: Apr. 8.208 UT, 9.2 (A. Pereira, Cabo

da Roca, Portugal; 10.121, 10.0 (K. Hornoch, Lelekovice, Czech Republic); 14.11, 9.6 (Hornoch); 21.16, 10.0 (Pereira); 23.13, 10.4 (P. Schmeer, Bischmisheim, Germany); 26.09, 10.3 (Hornoch); 27.15, 10.2 (J. M. Trigo, Benicassim, Castellon, Spain); 29.04, 10.4 (A. Baransky, Kiev, Ukraine).

1998 April 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green